What volume of dilute nitric acid of density 1.11g/mL and 19% nitric acid by weight contains in 10g nitric acid. 


1
Expert's answer
2021-04-16T05:00:11-0400

density D = m / V

V = m / D

V (HNO3) = 10 / 1.11 = 9 ml